If the productivity of s company is represented by the function

f(x,y)=40x3/4y1/4 with the utilisation of x units of the capital. Find

i) Fx and Fy and interpret

ii)If today the company is using 400 units and 250 units of labour and capital respectively.

determine the change in productivity if the labour is increased y 12 units and the capital is decreased by 8 units.


1
Expert's answer
2021-08-10T09:29:57-0400

"F(x,y)=40x^{3\/4}y^{1\/4}" - the productivity of a company,

"x" - number of units of the capital, "y" - number of units of labour

(i)

"F_x(x,y)=40\cdot \frac{3}{4}x^{3/4-1}y^{1/4}=30(y/x)^{1/4}"

"F_y(x,y)=40x^{3/4}\cdot \frac{1}{4}y^{1/4-1}=10(x/y)^{3/4}"

These formulae mean that the productivity of a company changes approximately with linear rate "F_x(x,y)=30(y\/x)^{1\/4}" per each new unit of the capital and "F_y(x,y)=10(x\/y)^{3\/4}" per each new unit of labour, when the current values of the capital and labour equal to "x" and "y" correspondently.

(ii)

If "x=400", "y=250" then

"F_x(x,y)=30(y\/x)^{1\/4}=30(250\/400)^{1\/4}=26.7"

"F_y(x,y)=10(x\/y)^{3\/4}=10(400\/250)^{3\/4}=14.2"

With these "x" and "y", if "\Delta y=12" and "\Delta x=-8", then the value of change will be approximately

"\Delta F\approx F_x\Delta x+F_y \Delta y=26.7\cdot(-8)+14.2\cdot12=-43.2"